OpenStudy (anonymous):

HELP! Determine whether the series is convergent or divergent. Tell what test that you used to make that determination.

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{ e^n }{ 3^{n-1}}\]

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{ \sqrt{n} }{ n-1 }\]